# Kinetic theory of point vortices at order ${1/N}$ and ${1/N^{2}}$

### v1

@@ II.3 Initial conditions
Figure 2: Illustration of the DF from Eq. (6) (full lines) and the considered frequency profiles (dashed lines): Profile (1) (top panel) is non-monotonic, while Profile (2) (bottom panel) is monotonic.
@@ II.4 Active fraction
To limit ourselves to the dynamically hot limit, i.e. the limit of small amplification of perturbations, we assume that part of the Hamiltonian is not self-consistently generated by the vortices. As such, we introduce an active fraction, ${0\!<\!q\!<\!1}$, and make the replacement
@@
In that expression, ${H_{\epsilon}[F](\mathbf{r})\!=\!\!\int\!\mathrm{d}\mathbf{r}^{\prime}F(\mathbf{r}^{\prime},t\!=\!0)\,U_{\epsilon}(\mathbf{r},\mathbf{r}^{\prime})}$ is the (softened) potential generated by the initial DF from Eq. (6). We also introduced ${H_{0}\!=\!H_{0}(J)}$ an axisymmetric external potential. Importantly, with the present choice, at ${t\!=\!0}$, we have ${\langle H_{\mathrm{tot}}\rangle\!=\!\sum_{i}\gamma_{i}\,H_{0}(J)}$, i.e. the initial mean-field orbital frequencies are simply ${\Omega(J)\!=\!\mathrm{d}H_{0}/\mathrm{d}J}$. From it, we also define the dynamical time as ${T_{\mathrm{dyn}}\!=\!2\pi/|\Omega(J_{0})|}$. With this setup, changing the value of $N$ amounts to changing the amplitude of the Poisson fluctuations, while changing $q$ amounts to changing the dynamical temperature, i.e. the level of self-consistent amplification in the system. Importantly, changing $N$ or changing $q$ does not lead to any change in the mean frequency profile. In the following, for simplicity, we assume that the point vortices have the same circulation, ${\gamma_{i}\!=\!\gamma}$.
@@ II.5 Frequency profiles
For the frequency profiles, we consider two different cases, profiles (1) and (2), as illustrated in Fig. 2. As detailed in Appendix A, these correspond to the (unsoftened) potential generated by a simple axisymmetric background distribution (Eq. 37). [^6] We emphasise that the orbital frequency (angular velocity) profile ${\Omega(J)}$ is not self-consistently generated by the point vortices that constitute our system, i.e. by the DF from Eq. (6), but is imposed from the outside, e.g., being generated by a background distribution of point vortices exterior to our system. This allows us to prescribe the orbital frequency in a convenient manner. The main difference between these two frequency profiles, ${J\!\mapsto\!\Omega(J)}$, is that profile (1) is non-monotonic while profile (2) is monotonic. As the upcoming section will detail, this plays a crucial role in defining the system’s regime of relaxation.
III ${1/N}$ dynamics
@@ III.1 Kinetic equation
First, let us consider profile (1). At order ${1/N}$, relaxation is driven by two-body effects. In the limit where collective effects can be neglected, this is described by the ${1/N}$ inhomogeneous Landau equation (see, e.g., Chavanis, 2013). In the present context, it reads
@@
[…]
@@
[…]
@@
the ${1/N}$ Landau equation satisfies an $H$-theorem, i.e. it drives ${\mathrm{d}S/\mathrm{d}t\!\geq\!0}$. Finally, the thermodynamical equilibria, i.e. the Boltzmann DFs, of the form
@@
are generic steady states of the ${1/N}$ Landau equation, i.e. they satisfy ${\partial F_{\mathrm{B}}/\partial t\!=\!0}$ when inserted into Eq. (10). There is no guarantee, however, that the inhomogeneous Landau equation at order ${1/N}$ relaxes towards the Boltzmann distribution because of (local) kinetic blockings, as explained below.
Performing explicitly the sum over $k$ in Eq. (10), the ${1/N}$ Landau equation can be rewritten as
@@
[…]
@@
the total coupling coefficient. In that form, Eq. (14) involves the simple resonance condition ${\delta_{\mathrm{D}}[\Omega(J)\!-\!\Omega(J_{1})]}$. If the frequency profile is monotonic, the condition ${\Omega(J_{1})\!=\!\Omega(J)}$ implies ${J_{1}\!=\!J}$ and Eq. (10) vanishes, i.e. ${\partial F/\partial t\!=\!0}$. There is no relaxation through ${1/N}$ effects. This is a situation of global kinetic blocking at order ${1/N}$. In the presence of a non-monotonic frequency profile, for a given $J$, there may exist another ${J_{1}\!\neq\!J}$ such that ${\Omega(J_{1})\!=\!\Omega(J)}$. In that case, Eq. (10) does not vanish and the system can relax through ${1/N}$ effects. This is the case for profile (1), as visible in Fig. 2.
@@
with respect to the total number of particles, $N$, and the system’s active fraction, $q$. As a result, it is useful to introduce the rescaled relaxation rate
$\mathcal{R}_{1}(J,t)=\frac{\partial N(\!<\!J,t)/\partial t}{N/T_{\mathrm{dyn}}}\,\frac{N}{q^{2}},$ (17)
with ${N(\!<\!J,t)\!=\!(2\pi/\gamma)\int_{0}^{J}\mathrm{d}J^{\prime}F(J^{\prime},t)}$ the number of particles with an action smaller than $J$ at time $t$. Its time derivative is given by ${\partial N(\!<\!J,t)/\partial t\!=\!(2\pi/\gamma)\mathcal{F}(J,t)}$, with the Landau flux ${\partial F(J,t)=\partial\mathcal{F}(J,t)/\partial J}$, following Eq. (14). Here, $\mathcal{R}_{1}$ is a dimensionless quantity, but more importantly, following Eq. (16), the ${1/N}$ Landau equation predicts that $\mathcal{R}_{1}$ should be independent of both $N$ and $q$.
@@ III.2 Linear stability
Let us now ensure that we consider systems that are dynamically hot enough. As detailed in Appendix C, the strength of collective effects can be estimated via the “dielectric matrix”, ${\mathbf{E}_{k}(\omega_{\mathrm{R}})}$ for ${\omega_{\mathrm{R}}\!\in\!\mathbb{R}}$. Then, the dressed coupling coefficients, ${\Lambda^{\mathrm{d}}_{k}}$, that appear in the ${1/N}$ Balescu–Lenard equation when compared with the ones from the Landau equation scale typically like ${\Lambda^{\mathrm{d}}_{k}(\mathbf{J};\omega_{\mathrm{R}})\!\sim\!\Lambda_{k}(\mathbf{J})/|\mathbf{E}_{k}(\omega_{\mathrm{R}})|}$, with ${|\mathbf{E}_{k}(\omega_{\mathrm{R}})|\!=\!\det[\mathbf{E}_{k}(\omega_{\mathrm{R}})]}$, the determinant of the dielectric matrix. In practice, the closer ${|\mathbf{E}_{k}(\omega_{\mathrm{R}})|}$ is from unity, the weaker are collective effects. On the contrary, if ${|\mathbf{E}_{k}(\omega_{\mathrm{R}})|\!=\!0}$, then the system supports an infinite amplification, i.e. this is a neutral mode. Importantly, since it describes collisionless response theory, we note that ${|\mathbf{E}_{k}(\omega_{\mathrm{R}})|}$ is independent of $N$ and depends only on the active fraction, $q$.
@@
Figure 3: Nyquist contour for profile (1) for various values of the active fraction $q$, for the resonance ${k\!=\!1}$. We refer to Appendix C for the definition of the “dielectric matrix”, ${\mathbf{E}_{k}(\omega)}$. For ${q\!\gtrsim\!0.0010}$, the system is found to be linearly unstable.
In that figure, we observe that for ${q\!\gtrsim\!0.0010}$, the system is linearly unstable since the Nyquist contour encloses the origin of the complex plane. [^7] See, e.g., Chavanis (2012d) for a simple illustration of the Nyquist method in homogeneous systems with long-range interactions. The possibly unstable profile (1) looks similar to the “gravity” case (compare Fig. 3 with fig. 2 of Chavanis (2012d)), while the always stable profile (2) looks similar to the “plasma” case (compare Fig. 10 with fig. 21 of Chavanis (2012d)). Moreover, we find that for ${q\!\lesssim\!0.0003}$, the contour remains close to ${|\mathbf{E}_{k}(\omega_{\mathrm{R}})|\!=\!1}$, i.e. collective amplification is small. In all the upcoming explorations, we limit ourselves to numerical simulations with ${q\!\leq\!0.0003}$, i.e. well within the range of applicability of the ${1/N}$ Landau equation in which collective effects are neglected.
@@ III.3 Small-scale contributions
Following Eq. (15), we can naturally decompose the relaxation rate from Eq. (17) as
$\mathcal{R}_{1}(J)\!=\!\sum_{k>0}\mathcal{R}_{1}(J,k),$ (18)
@@
where we dropped the dependence with respect to $t$ to shorten the notations. In Fig. 4, we illustrate the function ${k\!\mapsto\!\mathcal{R}_{1}(J\!=\!1.1\,J_{0},k)}$, emphasising that the considered action satisfies a non-local resonance condition (Fig. 2).
@@
Figure 4: Contribution to the relaxation rate, ${\mathcal{R}_{1}(k)}$ (Eq. 18), evaluated in ${J/J_{0}\!=\!1.1}$, as a function of the resonance number, $k$. The different thicknesses correspond to different values of the softening length, $\epsilon$. For small $k$, we recover the expected scaling in $k^{-1}$.
@@
From Fig. 4, we can make three observations: (i) For $k$ small enough, one finds ${\mathcal{R}_{1}(k)\!\propto\!k^{-1}}$, in agreement with the decay rate that can directly be inferred from Eq. (14), as detailed in Appendix D. (ii) As $k$ increases, we find that ${k\!\mapsto\!\mathcal{R}_{1}(k)}$ decreases faster than $k^{-1}$. This comes from the fact that for $k$ large enough, the difference between the two actions that are in resonance leads to a decay faster than algebraic (see, e.g., Eq. 45b). (iii) Finally, as one increases the softening length, $\epsilon$, (Eq. 4), the decay rate with respect to $k$ gets even faster. This makes perfect sense physically, since softening naturally washes out contributions from small scales. Overall, in the present regime, the ${1/N}$ Landau equation does not suffer from any small-scale divergence, contrary to the small-scale Coulomb logarithm that plagues the ${1/N}$ Landau equation when applied to ${3D}$ self-gravitating systems (see, e.g. Chavanis, 2013).
III.4 Relaxation rate

### v2

@@ II.3 Initial conditions
Figure 2: Illustration of the DF from Eq. (6) (full lines) and the considered frequency profiles (dashed lines). Profile (1) (top panel) is non-monotonic and has an extremum in $J_{\star}\!=\!1.03\,J_{0}$}, while profile (2) (bottom panel) is monotonic.
@@ II.4 Active fraction
To limit ourselves to the dynamically hot limit, i.e. the limit of small amplification of perturbations, we assume that part of the Hamiltonian is not self-consistently generated by the vortices. As such, we introduce an active fraction, 0<q<1{0\!<\!q\!<\!1}, and make the replacement
@@
In that expression, $H_{\epsilon}[F](\mathbf{r})\!=\!\!\int\!\mathrm{d}\mathbf{r}^{\prime}F(\mathbf{r}^{\prime},t\!=\!0)\,U_{\epsilon}(\mathbf{r},\mathbf{r}^{\prime})$} is the (softened) potential generated by the initial DF from Eq. (6). We also introduced $H_{0}\!=\!H_{0}(J)$} an axisymmetric external potential. Importantly, with the present choice, at t=0{t\!=\!0}, we have $\langle H_{\mathrm{tot}}\rangle\!=\!\sum_{i}\gamma_{i}\,H_{0}(J)$}, i.e. the initial mean-field orbital frequencies are simply $\Omega(J)\!=\!\mathrm{d}H_{0}/\mathrm{d}J$}. From it, we also define the dynamical time as $T_{\mathrm{dyn}}\!=\!2\pi/|\Omega(J_{0})|$}. With this setup, changing the value of $N$ amounts to changing the amplitude of the Poisson fluctuations, while changing $q$ amounts to changing the dynamical temperature, i.e. the level of self-consistent amplification in the system. Indeed, the smaller $q$, the less the point vortices contribute themselves to the overall total Hamiltonian. Importantly, changing $N$ or changing $q$ does not lead to any change in the mean frequency profile. In the following, for simplicity, we assume that the point vortices have the same circulation, $\gamma_{i}\!=\!\gamma$}.
@@ II.5 Frequency profiles
For the frequency profiles, we consider two different cases, profiles (1) and (2), as illustrated in Fig. 2. As detailed in Appendix A, these correspond to the (unsoftened) potential generated by a simple axisymmetric background distribution (Eq. 46). [^6] We emphasise that the orbital frequency (angular velocity) profile $\Omega(J)$} is not self-consistently generated by the point vortices that constitute our system, i.e. by the DF from Eq. (6), but is imposed from the outside, e.g., by being generated by a background distribution of point vortices exterior to our system. This allows us to prescribe the orbital frequency in a convenient manner. The main difference between these two frequency profiles, $J\!\mapsto\!\Omega(J)$}, is that profile (1) is non-monotonic, exhibiting a minimum in $J_{\star}\!\simeq\!1.03\,J_{0}$}, while profile (2) is monotonic. As the upcoming section will detail, this difference plays a crucial role in defining the system’s regime of relaxation.
III 1/N{1/N} dynamics
@@ III.1 Kinetic equation
First, let us consider profile (1). At order 1/N{1/N}, relaxation is driven by two-body effects. In the limit where collective effects can be neglected, this is described by the 1/N{1/N} inhomogeneous Landau equation (see, e.g., Chavanis, 2013). In the present context, it reads
@@
$\times{}$ $\,\delta_{\mathrm{D}}\big(\mathbf{k}\!\cdot\!\mathbf{\Omega}\big)\,\mathbf{k}\!\cdot\!\frac{\partial}{\partial\mathbf{J}}F_{2}(\mathbf{J})\bigg],$ (10)
where we dropped the time dependence for clarity. We refer to Appendix D for the expression of the coupling coefficients, $|\Lambda_{k}(\mathbf{J})|^{2}$}, which are independent of both $N$ and $q$. In Eq. (10), notations are shortened with the 2-vectors $\mathbf{J}\!=\!(J,J_{1})$}, $\mathbf{\Omega}\!=\!(\Omega[J],\Omega[J_{1}])$}, and $\mathbf{k}\!=\!(k,-k)$}, as well as $F_{2}(\mathbf{J})\!=\!F(J)F(J_{1})$}. When accounting for collective effects, Eq. (10) becomes the inhomogeneous Balescu–Lenard equation (Heyvaerts, 2010; Chavanis, 2012c). The functional form of Eq. (10) remains the exact same, and one only has to replace the bare coupling coefficients, $|\Lambda_{k}(\mathbf{J})|^{2}$}, with their (frequency-dependent) dressed analogs, $|\Lambda^{\mathrm{d}}_{k}(\mathbf{J};k\Omega(J))|^{2}$}. In the dynamically hot limit, as expected, one has $|\Lambda^{\mathrm{d}}_{k}|^{2}\!\to\!|\Lambda_{k}|^{2}$}. This is detailed in Appendix C, where we present a new approach to compute $|\Lambda^{\mathrm{d}}_{k}|^{2}$}.
@@
$\Gamma{}$ $=\!\!\int\!\!\mathrm{d}J\,F(J,t){}$ $\text{(total circulation)};$ (11a)
$L{}$ $=\!\!\int\!\!\mathrm{d}J\,J\,F(J,t){}$ $\text{(total momentum)};$ (11b)
$E{}$ $=\!\!\int\!\!\mathrm{d}J\,H(J)\,F(J,t){}$ $\text{(total energy)}.$ (11c)
@@
the 1/N{1/N} Landau equation satisfies an $H$-theorem, i.e. it drives $\mathrm{d}S/\mathrm{d}t\!\geq\!0$}. Finally, the thermodynamical equilibria, i.e. the Boltzmann DFs, of the form
@@
are generic steady states of the 1/N{1/N} Landau equation, i.e. they satisfy $\partial F_{\mathrm{B}}/\partial t\!=\!0$} when inserted into Eq. (10). There is no guarantee, however, that the inhomogeneous Landau equation at order 1/N{1/N} relaxes towards the Boltzmann distribution because of (local) kinetic blockings, as explained below.
Performing explicitly the sum over $k$ in Eq. (10), the 1/N{1/N} Landau equation can be rewritten as
@@
$\times{}$ $\,\big|\Lambda_{\mathrm{tot}}(J,J_{1})\big|^{2}\,\bigg(F(J_{1})\,\frac{\partial F}{\partial J}\!-\!F(J)\,\frac{\partial F}{\partial J_{1}}\bigg)\bigg],$
@@
the total coupling coefficient. In that form, Eq. (14) involves the simple resonance condition $\delta_{\mathrm{D}}[\Omega(J)\!-\!\Omega(J_{1})]$}.
Equation (14) can also be rewritten under the form of a Fokker–Planck equation (see, e.g. Chavanis, 2012c), via
$\frac{\partial F(J,t)}{\partial t}\!=\!-\frac{\partial}{\partial J}\bigg[\!D_{1}(J)\,F(J)\!\bigg]\!+\!\frac{1}{2}\frac{\partial^{2}}{\partial J^{2}}\bigg[\!D_{2}(J)\,\frac{\partial F}{\partial J}\!\bigg].$ (16)
Here, $D_{1}$ [resp. $D_{2}$] is the first-order [resp. second-order] diffusion coefficient in action space. In the following, we will focus our interest only on the second-order diffusion coefficient. It generically reads
$D_{2}(J)=2\gamma\!\!\int\!\!\mathrm{d}J_{1}\,\delta_{\mathrm{D}}\big[\Omega(J)\!-\!\Omega(J_{1})\big]\,\big|\Lambda_{\mathrm{tot}}(J,J_{1})\big|^{2}\,F(J_{1}).$ (17)
Physically, $D_{2}(J)$} describes the diffusion induced by stochastic potential fluctuations. For example, let us consider a zero-mass test particle, of initial action $J$, embedded in the present system,. Then, on long timescales, the growth of the variance of the action increment for this particle, i.e. $\Delta J(T)\!=\!J(T)\!-\!J(0)$}, is given by
$D_{2}(J)=\lim\limits_{T\to+\infty}\big\langle\big[\Delta J(T)\big]^{2}\big\rangle/T.$ (18)
Here, the ensemble average is taken over both the realisations of the background system and the initial phase of the test particle.
Coming back to Eq. (14) we note that if the frequency profile is monotonic, the condition $\Omega(J_{1})\!=\!\Omega(J)$} implies $J_{1}\!=\!J$} and Eq. (10) vanishes, i.e. $\partial F/\partial t\!=\!0$}. There is no relaxation through 1/N{1/N} effects. This is a situation of global kinetic blocking at order 1/N{1/N}. We note that such a vanishing does not occur for the diffusion coefficient from Eq. (17). In the presence of a non-monotonic frequency profile, for a given $J$, there may exist another $J_{1}\!\neq\!J$} such that $\Omega(J_{1})\!=\!\Omega(J)$}. In that case, Eq. (10) does not vanish and the system can relax through 1/N{1/N} effects. This is the case for profile (1), as visible in Fig. 2.
@@
with respect to the total number of particles, $N$, and the system’s active fraction, $q$. As a result, it is useful to introduce the rescaled relaxation rate and diffusion coefficients
$\mathcal{R}_{1}(J,t){}$ $=\frac{\partial N(\!<\!J,t)/\partial t}{N/T_{\mathrm{dyn}}}\,\frac{N}{q^{2}},$ (20a)
$\mathcal{D}(J){}$ $=\frac{D_{2}(J)}{J_{0}^{2}/T_{\mathrm{dyn}}}\,\frac{N}{q^{2}},$ (20b)
with $N(\!<\!J,t)\!=\!(2\pi/\gamma)\int_{0}^{J}\mathrm{d}J^{\prime}F(J^{\prime},t)$} the number of particles with an action smaller than $J$ at time $t$. Its time derivative is given by $\partial N(\!<\!J,t)/\partial t\!=\!(2\pi/\gamma)\mathcal{F}(J,t)$}, with the Landau flux $\partial F(J,t)\partial t\!=\!\partial\mathcal{F}(J,t)/\partial J$}, following Eq. (14). Here, $\mathcal{R}_{1}(J)$} and $\mathcal{D}(J)$} are both dimensionless quantities. More importantly, the 1/N{1/N} Landau equation (Eq. 14) predicts that the two of them should be independent of both $N$ and $q$.
@@ III.2 Linear stability
Let us now ensure that we consider systems that are dynamically hot enough. As detailed in Appendix C, the strength of collective effects can be estimated via the “dielectric matrix”, $\mathbf{E}_{k}(\omega_{\mathrm{R}})$} for $\omega_{\mathrm{R}}\!\in\!\mathbb{R}$}. Then, the dressed coupling coefficients, $\Lambda^{\mathrm{d}}_{k}$}, that appear in the 1/N{1/N} Balescu–Lenard equation when compared with the ones from the Landau equation scale typically like $\Lambda^{\mathrm{d}}_{k}(\mathbf{J};\omega_{\mathrm{R}})\!\sim\!\Lambda_{k}(\mathbf{J})/|\mathbf{E}_{k}(\omega_{\mathrm{R}})|$}, with $|\mathbf{E}_{k}(\omega_{\mathrm{R}})|\!=\!\det[\mathbf{E}_{k}(\omega_{\mathrm{R}})]$}, the determinant of the dielectric matrix. In practice, the closer $|\mathbf{E}_{k}(\omega_{\mathrm{R}})|$} is from unity, the weaker are collective effects. On the contrary, if $|\mathbf{E}_{k}(\omega_{\mathrm{R}})|\!=\!0$}, then the system supports an infinite amplification, i.e. this is a neutral mode. Importantly, since it describes collisionless response theory, we note that $|\mathbf{E}_{k}(\omega_{\mathrm{R}})|$} is independent of $N$ and depends only on the active fraction, $q$.
@@
Figure 3: Nyquist contour for profile (1) for various values of the active fraction $q$, for the resonance k=1{k\!=\!1}. We refer to Appendix C for the definition of the “dielectric matrix”, $\mathbf{E}_{k}(\omega)$}. For $q\!\gtrsim\!0.0010$}, the system is found to be linearly unstable.
In that figure, we observe that for $q\!\gtrsim\!0.0010$}, the system is linearly unstable since the Nyquist contour encloses the origin of the complex plane. [^7] See, e.g., Chavanis (2012d) for a simple illustration of the Nyquist method in homogeneous systems with long-range interactions. The possibly unstable profile (1) looks similar to the “gravity” case (compare Fig. 3 with fig. 2 of Chavanis (2012d)), while the always stable profile (2) looks similar to the “plasma” case (compare Fig. 11 with fig. 21 of Chavanis (2012d)). Moreover, we find that for $q\!\lesssim\!0.0003$}, the contour remains close to $|\mathbf{E}_{k}(\omega_{\mathrm{R}})|\!=\!1$}, i.e. collective amplification is small. In all the upcoming explorations, we limit ourselves to numerical simulations with $q\!\leq\!0.0003$}, i.e. well within the range of applicability of the 1/N{1/N} Landau equation in which collective effects are neglected.
@@ III.3 Small-scale contributions
Following Eq. (15), we can naturally decompose the relaxation rate and diffusion coefficient from Eq. (20) as
$\mathcal{R}_{1}(J){}$ $=\sum_{k>0}\mathcal{R}_{1}(J,k),$ (21a)
$\mathcal{D}(J){}$ $=\sum_{k>0}\mathcal{D}(J,k),$ (21b)
@@
where we dropped the dependence with respect to $t$ to shorten the notations. In Fig. 4, we illustrate the dependence with $k$ of $\mathcal{R}_{1}(J,k)$} and $\mathcal{D}(J,k)$}, when evaluated in $J\!=\!1.1\,J_{0}$}.
@@
Figure 4: Contribution to the relaxation rate, $\mathcal{R}_{1}(J,k)$} (Eq. 21a, top panel), and diffusion coefficient, $\mathcal{D}(J,k)$} (Eq. 21b, bottom panel), evaluated in $J/J_{0}\!=\!1.1$}, as a function of the resonance number, $k$. The different colours correspond to different values of the softening length, $\epsilon$. For small $k$, we recover the expected scaling in $k^{-1}$, before it being damped by softening.
We emphasise that the considered action satisfies a non-local resonance condition (Fig. 2).
@@
From Fig. 4, we can make three observations. (i) For $k$ small enough, one finds $\mathcal{R}_{1}(J,k),\mathcal{D}(J,k)\!\propto\!k^{-1}$}, in agreement with the decay rate of the coupling coefficient, $k\!\mapsto\!U_{k}(J,J_{1})$}, that can directly be inferred from Eq. (52b), as detailed in Appendix D. (ii) As $k$ increases, we find that $k\!\mapsto\!\mathcal{R}_{1}(J,k)$} decreases faster than $k^{-1}$. This comes from the fact that for $k$ large enough, the difference between the two actions, $J$ and $J_{1}$, that are in resonance in Eq. (14) leads to a decay faster than algebraic (see Eq. 52b). Indeed, following Eq. (52b), and assuming $J_{1}\!\leq\!J$}, we find $U_{k}(J,J_{1})\!\propto\!(J_{1}/J)^{k}/|k|$}, hence driving an exponentially fast decay. (iii) On the contrary, as $k$ increases, $k\!\mapsto\!\mathcal{D}(J,k)$} continues to scale like $k^{-1}$. This is because local resonances, i.e. $J_{1}\!=\!J$} in Eq. (17), always contribute to the diffusion coefficient. Indeed, following Eq. (52b), for local resonances, one finds $U_{k}(J,J)\!\propto\!1/|k|$}. (iv) Finally, as one increases the softening length, $\epsilon$ (Eq. 4), the decay rate with respect to $k$ gets faster, for both the relaxation rate and the diffusion coefficient. Indeed, even when evaluated for local resonances, Eq. (54b) roughly gives the scaling $U_{k}(J,J)\!\propto\!(1\!-\!\epsilon/J)^{k}/|k|$}, hence imposing an exponentially fast decay. This makes perfect sense physically, since softening naturally washes out contributions from small scales.
Overall, in the present regime, the 1/N{1/N} Landau diffusion coefficient without softening suffers from a small-scale logarithmic divergence. This is similar to the Coulomb logarithm that plagues the 1/N{1/N} Landau equation when applied to 3D{3D} self-gravitating systems (see, e.g. Chavanis, 2013). In the presence of softening, this small-scale divergence of $\mathcal{D}(J,k)$} is naturally washed out. By contrast, the 1/N{1/N} Landau flux, because it is unaffected by local resonances, does not suffer from any small-scale divergence, even without softening. We also refer to Appendix H for a discussion of how resonance broadening also cures this small-scale divergence. In practice, in all the upcoming simulations, we will consider dynamics driven by the softened interaction potential from Eq. (4), fixing the softening length to $\epsilon\!=\!10^{-2}J_{0}$}.
III.4 Relaxation and diffusion